OpenStudy (anonymous):

brianna has 11 2/3 cups of milk ` she will certainly need 1 1/6 cups of milk to make a pot of hot cocoa ; how many pots of hot cocoa can she make ?

OpenStudy (anonymous):

you have to divide for this one \[11\tfrac{2}{3}\div1\tfrac{1}{6}\]

OpenStudy (anonymous):

again turn each improper fraction in to a mixed number, the to divide invert and multiply

OpenStudy (anonymous):

11 2/3 / 1 1/6 = 35/3 / 7/6 = 35/3 * 6/7 = 210/21 = 10

OpenStudy (anonymous):

\[11\frac{2}{3}=\frac{35}{3}\] \[1\tfrac{1}{6}=\frac{7}{6}\]